# Contextual Prompt Learning for Vision-Language Understanding

### v1

@@ Abstract
Recent advances in multimodal learning has resulted in powerful vision-language models, whose representations are generalizable across a variety of downstream tasks. Recently, their generalizability has been further extended by incorporating trainable prompts, borrowed from the natural language processing literature. While such prompt learning techniques have shown impressive results, we identify that these prompts are trained based on global image features which limits itself in two aspects: First, by using global features, these prompts could be focusing less on the discriminative foreground image, resulting in poor generalization to various out-of-distribution test cases. Second, existing work weights all prompts equally whereas our intuition is that these prompts are more specific to the type of the image. We address these issues with as part of our proposed Contextual Prompt Learning (CoPL) framework, capable of aligning the prompts to the localized features of the image.
Our key innovations over earlier works include using local image features as part of the prompt learning process, and more crucially, learning to weight these prompts based on local features that are appropriate for the task at hand. This gives us dynamic prompts that are both aligned to local image features as well as aware of local contextual relationships. Our extensive set of experiments on a variety of standard and few-shot datasets show that our method produces substantially improved performance when compared to the current state of the art methods. We also demonstrate both few-shot and out-of-distribution performance to establish the utility of learning dynamic prompts that are aligned to local image features.
1 Introduction
Figure 1: Left hand side picture describes the arechitecure of the CoCoOp and right hand side highlights the architecture of CoPL. In case of CoCoOp, the global image features considered as the coditioned vector for the prompt tokens where as for our CoPL model, we consider localized image features. The colour scheme highlights the weighted prompts according to their semantic alignment with the local image features.
@@ 1 Introduction
Fully supervised computer vision models for problems like classification are typically trained on datasets like ImageNet [5], OpenImages [32], JFT300M [29, 57] etc., and have also proven themselves to be effective for a variety of downstream tasks via transfer learning [50, 20, 59]. Despite this, it is challenging to adapt these models to other domains due to various reasons including limited data, annotation constraints etc. Additionally, since these models are trained for specific objectives like classification, they tend to capture concepts related to categories seen during training and not to scale to unseen classes during inference.
@@
To address the finetuning issue above, there has been recent effort in tuning the associated prompts (instead of the model weights). Inspired by traditional prompt engineering efforts [49, 3, 53] , there has been some work in tuning discrete prompts from predefined prompt templates [51] that help in capturing rich semantics from user intents and align them to visual contents. However, since building a rich semantic based prompt templates require domain specific and linguistic knowledge, it is tedious and time consuming job to perform. The CoOp [63] algorithm used ideas from soft-prompting in natural language processing [17, 26, 33, 37, 43, 42] to train dynamic learnable prompt vectors with backpropagation and preserve the semantic relationship between sentences and labels [43]. However, the context lerned with CoOp in this fashion fails to generalize to unseen classes, leading to the need for dynamically updating prompts based on the image context, an idea that was proposed in CoCoOp [62]. This model was trained by explicitly conditioning the prompts on image feature vectors as tokens where a separate lightweight neural network (called meta-net in their work) was used to equally weight all the prompt vectors.
@@
In our work, we identify two key issues with the aforementioned architecture. First, the features obtained from meta-net in CoCoOp [62] are global in nature and hence susceptible to issues like clutter and noise in many few-shot and out-of-distribution test cases (we demonstrate this empirically later on). Next, these features are directly added to all the learned prompt vectors, thus resulting in an equal weighting for each prompt. Consequently, this model is unable to learn which of the prompt vectors are more semantically relevant and contextually meaningful during inference, which makes model less generalizable on unseen classes in a zero-shot settings.
@@
To address the aforementioned problems, we propose a new technique called Contextual Prompt Learning (CoPL). Our key ideas include aligning prompts to local image context, realized with local features, and determining which prompts are more sematically relevant conditioned on such local context. Our insight is that by doing so, we are able to learn a more appropriate weighting of the prompts, unlike equal weighting of prior work above, that is semantically reflective of the actual content of the image under consideration. Inspired by the concept of global attention from the NLP literature [44], during training, we propose to align each local feature vector (e.g., computed from a local image patch) to a set of dynamic soft-prompts using a learned context vector that attends to these prompt vectors. An overview of the methdology is shown in Figure 1. This produces a set of attention weights for the prompt vectors that are semantically aligned to local image regions. This results in CoPL learning more generalizable features as we demonstrate with an extensive set of zero- and few-shot classification evaluation results.
@@
We conduct a comprehensive set of experiments on visual classification on $11$ different datasets (see Section 4 and scenarios (zero-shot, one-shot, seen/unseen, and within-dataset and cross-dataset). Across all these experiments, we demonstrate substantial performance improvement when compared to our closest baselines CoOp [63] and CoCoOp [62], indicating the ability of our method to be adapted across various classification settings with little or no training and much reduced prompt engineering.
Our key contributions are summarized below:
  • We identify two key issues with existing prompt-based image classification methods: equal weighting of the prompt vectors and no flow of contextual local information of input images to the prompt vectors while learning during back-propagation.
  • We propose CoPL: Contextualized Prompt Learning, a new method that addresses the issues above by learning prompt weights dynamically and aligning the resulting prompt vectors with local image features.
@@
  • We conduct extensive experiments with CoPL under a variety of classification scenarios and demonstrate substantial performance improvements, in particular in various unseen and few-shot data scenarios. Very interesting to see that, on $11$ different image recognition dataset, on average CoPL achieves state-of-the-art performance on unseen class classifications beating state-of-the-art zeros-hot large-scale model CLIP by $1.4\%$ and conditional prompting model CoCoOp by $3.9\%$ on accuracy. We evaluated CoPL on cross-dataset zero-shot image recognition tasks and on an average on $8$ datasets, CoPL outperformed CoCoOp by $2.3\%$ on accuracy.
2 Related Work
Multimodal Models
@@ 2 Related Work
Vision-language models have shown great potential in learning generic visual representations. The core idea has been to use natural language supervision for image representation learning and align them jointly in the same embedding space [25, 51]. A vision-language model can be seen as a composition of three parts - a text encoder, an image encoder and a learning methodology that leverages information from both these modalities. Early explorations in this line of research included the problem formulations - metric learning [13], multilabel classification [18, 27], n-gram language learning [34], captioning [8]. Traditionally, hand-crafted descriptors [11, 55] were the mode of capturing image representations. With the introduction deep neural nets, researchers used convolutional neural net based architectures [14, 1] to image representations where as the texts are encoded by simply taking TF-IDF features of the words. With the introduction of pre-trained word vectors the representations learnt by capturing semantic relativity of the words improved [14, 55].
Recent works have focused on learning joint representations of both the modalities using deep learning architectures [15, 25, 38, 28]. As mentioned in [63], there have been significant developments in the areas of transformers[58], and contrastive representation learning [24, 21]. Li et al., [35] proposed VISUALBERT, where texts and images are jointly encoded in a single transformers architecture. The core idea was to align both the modalities inside the transformers architecture by leveraging the cross-attention mapping. The text embddings are obtained from the BERT language model [9] and the image embeddings are obtained from a carefully designed image encoder. One of the biggest release in the multimodal space is the CLIP model [51]. It is a dual encoder based model and during training it matches pairs of images and texts. It is trained on $400$ million training dataset and achieved state-of-the-art performance on different zero-shot downstream tasks. One of the main component of CLIP is the carefully designed prompts which most of the times very hard to formulate. To overcome this, Zhou et al., [63] designed CoOp, which trains dynamic soft-prompts during back-propagation. They keep the underneath multimodal models fixed and only trains the prompt tokens. CoOp has produced impressive results on different image recognition tasks.
Prompting
The intuition behind prompt learning is to capture user intention and instructions to perform certain downstream tasks. In natural language processing literature, capturing instructions to perform various downstream tasks is well established [39, 48, 64]. With the introduction of GPT [3], prompt engineering is shown to be performing efficiently in few-shot knowledge adaptation. Recently Liu et al., [40] designed a template-based question answering model to perform downstream named entity recognition tasks. In their work, they referred to the classical template prompting setup where the templates consist of questions as prompts “What is the [E]?”. In a probabilistic setup, the model predicts the “[E]” token from the distribution of classes which completes the sentence. For example: if the class is “location”, then the model predicts the token and fills in the “[E]” to complete the sentence as “What is the location?”. The key problem behind building prompt templates is to identify the format and linguistic relations with the classes within the dataset. Thus, it requires high skills and knowledge to come up with good prompt templates [41]. Jiang et al., [26] proposed candidate prompts can be generated using text mining, relying on the knowledge of the language model. Deng et al., [7] came up with solution to optimize the discrete prompts using reinforcement learning. The architecture consists of parameter-efficient policy network which generates the discrete prompts after optimizing. They have implemented effective reward stabilization process to tune the rewards efficiently using language models.
Recently, researchers have proposed “soft-prompting” which overcome the shortcomings of the template based prompting including the carefully prompt designing and engineering. The main intuition is to learn dynamic continuous prompt tokens during back-propagation [17, 26, 33, 37, 43, 42]. During optimization, the weights of the language models are fixed. During training, the models predict the classes, carefully mapped from the vocabulary called verbalizers. Recently Goswami et al., [19] in their work highlights the soft prompts can be further tuned with the semantic knowledge of the language models for the different domain adaptation without explicitly verbalizers setup.
Fine-tuning
Before the adaptation to prompt based learning, the standard process was to fine-tune the existing pretrained models on the downstream vision tasks [31, 54, 22]. Gao et al., proposed CLIP-Adapter [16] to improve efficiency of the CLIP by applying a residual transformation layer over the feature embedding or classifier weight. Wortsman et al., [60] proposed WiSE-FT which fine-tunes the layers as a post-ensemble method to improve the efficacy of CLIP.
3 Methodology
Figure 2: The architecture of CoPL: Contextualized Prompt Learning
@@ 3 Methodology
Here, we introduce and discuss details of our proposed method we call CoPL: Contextualized Prompt Learning. Since CoPL uses the same architectural backbone as CLIP [51], we first begin with a brief review followed by a discussion on how our closest baseline, CoCoOp [62], is trained. This will help set the stage for motivating and discussing our method which we do subsequently. While we use the CLIP backbone for simplicity, we are in no way limited by this design choice. In fact, our method is very much applicable to be used in conjunction with a variety of other architectures, e.g., VisualBERT [35], MDETR [28], GLIP [36] etc.
3.1 Review of CLIP
@@
CLIP [51] is trained using the standard contrastive learning setup where there is a text encoder and an image encoder and the overall objective function is to get their outputs as close as possible in the joint space. The text encoder is implemented with the Transformer model [58] which takes word sequences as input and produces both individual sequence-level as well as overall sentence-level representations. The image encoder is implemented with the ViT architecture [10] which produces local (at the patch level) as well as global image features. The contrastive loss function is designed to capture the similarity between the relevant text and images, that is, the cosine similarity between the related text and image will be maximized whereas the cosine similarity with all other unrelated pairs will be minimized. To be precise, for a $K$-class classification problem, the discrete prompts are designed to have one “class” token. The training objective is then to fill the token with the $i$-th class name where $w_{i}$ is the weight vector generated by the text-encoder for the same.
$p(y|x)=\frac{exp(sim(x,w_{y}))/\gamma}{\sum_{i=1}^{k}exp(sim(x,w_{i}))/\gamma}$ (1)
where $x$ is the image feature vector produced by the image encoder and $\gamma$ is a temperature parameter.
3.2 Review of CoCoOp
@@
CoCoOp: Conditional Context Optimization is built on top of a previously published Context Optimization (CoOp) [63] algorithm. It is observed that the choice of prompts plays a major role in vision-language understanding [51]. However, finding the right match between prompts and description of the image often takes a significant amount of time [63] and sometimes can change the model results drastically. For instance, adding task-specific manual prompts in the Flowers102 dataset [46] would include the word “flowers”, whereas for the DTD dataset [4] the word should be “texture”. In both cases, adding these words in prompts requires manual intervention and careful observation. To overcome such a need for prompt engineering, the CoOp model [63] trains a set of continuous vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}\}$ as context tokens during back-propagation. These vectors are represented as word vectors instead of discrete predefined prompts. Specifically, the prompts for the $i$-th class, denoted by $t_{i}$ can be represented as $[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M},\mathbf{c}_{i}]$, where $M$ is the number of learnable prompt vectors and $\mathbf{c}_{i}$ is the class embeddings. The CoCoOp [62] algorithm improves CoOp’s [63] performance learning to generate prompts conditioned on each image instance, i.e., these now change for each image unlike in CoOp where they are fixed. CoCoOp does this by training an additional two-layer neural network called meta-net that takes an image feature vector as input and produces a conditional vector that is combined with the prompt vectors to generate the final image-dependent prompts. This is done as follows:
$\mathbf{v}_{m}(\mathbf{x})=\mathbf{v}_{m}+\mathbf{\pi}$ (2)
$\mathbf{\pi}=h_{\theta}(\mathbf{x})$ (3)
@@
3.3 CoPL: Contextual Prompt Learning
While CoCoOp generalizes better than CoOp for unseen-class classification, many issues remain. First, since CoCoOp uses global feature vectors for learning the updated prompts, it focuses less on the discriminative regions in images (which tend to be more local). Next, the addition operation performed in Equation 2 does not capture the individual importance of each prompt token. This is particularly important since certain discriminative regions in images may weight certain prompts more when compared to others and this is not captured in the CoCoOp model. To address these issues, we propose CoPL, a simple and intuitive algorithm that operates at the local feature granularity while also aligning them with prompts to learn prompt importance weights.
Given an image $\mathbf{I}$, we first compute a set of local feature vectors. For instance, this can be the output of a vision transformer model that generates patch embeddings. Let $\mathbf{s}\in\mathbb{R}^{P\times B\times d}$ (where $P$ is number of patches from image, $B$ is training batch size, and $d$ is feature dimesionality) be this. Conditioned on these local features, we determine the semantically most meaningful prompts. To do this, we take inspiration from the attention work of Luong et al. [44] and generate context representations that explicitly consider both the learnable prompt tokens as well as the patch representations. We first learn a lightweight neural network to generate a conditional token for the representation of each patch in $\mathbf{s}$ as:
@@ 3.3 CoPL: Contextual Prompt Learning
$=\frac{\text{exp}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i}))}{\sum_{i=1}^{M}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i}))}$
@@
$\mathbf{v}_{m}(\mathbf{x})=\sum_{i=1}^{P}\mathbf{v}_{m}+\mathbf{c}_{i}$ (8)
In a nutshell CoPL calculates the prompts for the $i$-th class as $t_{i}=[v_{1}(x),v_{2}(x),...,v_{M}(x),cl_{i}]$, where $cl_{i}$ is the $i$-th class. The prediction probability is calculated as
@@
where $g()$ is the feature vector produced by the text encoder. During our entire pipeline, the pre-trained CLIP model is fixed.
4 Experimental Evaluation
Datasets ($\downarrow$) Protocols ($\downarrow$) CLIP CoOp CoCoOp CoPL
Caltech101 Seen 96.8 98.0 97.9 98.1
Unseen 94.0 89.8 93.8 94.9
HM 95.4 93.7 95.8 96.5
OxfordPets Seen 91.1 93.6 95.2 95.6
Unseen 97.2 95.2 97.6 97.8
HM 94.1 94.4 96.4 96.7
StandfordCars Seen 62.3 78.1 70.4 70.7
Unseen 74.8 60.4 73.5 74.4
HM 68.6 68.1 72.0 72.6
Flowers102 Seen 72.0 97.6 94.8 96.1
Unseen 77.8 59.6 71.7 72.1
HM 74.8 74.0 81.7 84.1
Food101 Seen 90.1 88.3 90.7 90.9
Unseen 91.2 82.2 91.2 91.4
HM 90.6 85.1 90.9 91.1
FGVCAircraft Seen 27.1 40.4 33.4 36.1
Unseen 36.2 22.3 23.7 31.3
HM 31.0 28.7 27.7 33.7
SUN397 Seen 69.3 80.6 79.7 80.2
Unseen 75.3 65.8 76.8 77.3
HM 72.2 72.5 78.2 78.7
DTD Seen 53.2 79.4 77.0 78.0
Unseen 59.9 41.1 56.0 50.0
HM 56.3 54.2 64.8 64.0
ImageNet Seen 72.4 76.4 75.9 77.8
Unseen 68.1 67.8 70.4 71.3
HM 70.2 71.9 73.1 74.5
EuroSAT Seen 56.4 92.1 87.4 89.2
Unseen 64.0 54.7 60.0 34.2
HM 60.0 68.9 71.2 61.7
UFC101 Seen 70.5 84.6 82.3 83.1
Unseen 77.5 56.0 73.4 76.6
HM 73.8 67.4 77.6 79.8
Average Seen 69.3 82.7 80.5 81.4
Unseen 74.22 63.2 71.7 75.6
HM 71.7 71.6 75.8 78.5
Table 1: Performance of CLIP, CoOP, CoCoOp, and CoPL on $11$ classification datasets; each of the training dataset consists $16$-shots per class. The dataset highlights CoPL do better generalization than other methods. The baseline results are taken from CoCoOp paper [62]. HM refers to Harmonic Mean between the seen and unseen classes.
Our key insight in CoPL is that we are able to learn generalizable representations by focusing more on the local representations of the image, along with deferentially weighting the learned prompt embeddings. To bring out the efficacy of this, we evaluate our method on multiple settings: $(i)$ knowledge transfer to unseen classes from seen classes within the same dataset, $(ii)$ zero-shot performance across multiple datasets, $(iii)$ efficiency on unseen classes while trained with $1$-shot training set within the same dataset.

### v2

# CoPL: Contextual Prompt Learning for Vision-Language Understanding

@@ Abstract
Recent advances in multimodal learning has resulted in powerful vision-language models, whose representations are generalizable across a variety of downstream tasks. Recently, their generalization ability has been further extended by incorporating trainable prompts, borrowed from the natural language processing literature. While such prompt learning techniques have shown impressive results, we identify that these prompts are trained based on global image features which limits itself in two aspects: First, by using global features, these prompts could be focusing less on the discriminative foreground image, resulting in poor generalization to various out-of-distribution test cases. Second, existing work weights all prompts equally whereas intuitively, prompts should be reweighed according to the semantics of the image. We address these as part of our proposed Contextual Prompt Learning (CoPL) framework, capable of aligning the prompts to the localized features of the image. Our key innovations over earlier works include using local image features as part of the prompt learning process, and more crucially, learning to weight these prompts based on local features that are appropriate for the task at hand. This gives us dynamic prompts that are both aligned to local image features as well as aware of local contextual relationships. Our extensive set of experiments on a variety of standard and few-shot datasets show that our method produces substantially improved performance when compared to the current state of the art methods. We also demonstrate both few-shot and out-of-distribution performance to establish the utility of learning dynamic prompts that are aligned to local image features.
Introduction
Figure 1: The figure summarises our methodological contribution. CoPL leverages the local information in the image patches to learn contextualized prompts (left). Further, it weights each of the prompts differently, based on its semantic affinity. The coloured patches highlights how the prompts are weighted according to their semantic alignment with the local image features. Baselines approaches (on right) uses global image features for prompt learning, and weights each of the prompts equally.
@@ Introduction
Fully supervised computer vision models for problems like classification are typically trained on datasets like ImageNet (Deng et al. 2009a), OpenImages (Kuznetsova et al. 2018), JFT300M (Kolesnikov et al. 2020; Sun et al. 2017) etc., and have also proven themselves to be effective for a variety of downstream tasks via transfer learning (Plested and Gedeon 2022; Guo et al. 2019; Wan et al. 2019). Despite this, it is challenging to adapt these models to other domains due to various reasons including limited data and annotation overhead. Additionally, since these models are trained for specific objectives like classification, they tend to capture concepts related to categories seen during training and not to scale to unseen classes during inference.
@@
To enhance the adaptability of such models, there has been recent efforts in tuning the associated prompts (instead of the model weights). Inspired by traditional prompt engineering efforts (Petroni et al. 2019; Brown et al. 2020; Schick and Schütze 2021), there has been work in tuning discrete prompts from predefined prompt templates (Radford et al. 2021) that help in capturing rich semantics from user intents and align them to visual contents. However, since building a rich semantic based prompt templates require domain specific and linguistic knowledge, this apporach is not scalable. The CoOp (Zhou et al. 2022b) algorithm used ideas from soft-prompting in natural language processing (Gao, Fisch, and Chen 2021; Jiang et al. 2020; Lester, Al-Rfou, and Constant 2021; Li and Liang 2021; Liu et al. 2021, 2022) to train dynamic learnable prompt vectors with backpropagation and preserve the semantic relationship between sentences and labels (Liu et al. 2021). However, the context learned with CoOp in this fashion fails to generalize to unseen classes, leading to the need for dynamically updating prompts based on the image context, an idea that was proposed in CoCoOp (Zhou et al. 2022a). This model was trained by explicitly conditioning the prompts on image feature vectors as tokens where a separate lightweight neural network (called meta-net in their work) was used to equally weight all the prompt vectors. Later Yao, Zhang, and Xu (2023) note that CoOp-based methods suffer from catastrophic knowledge forgetting, where these methods gradually miss out on essential general textual knowledge. This leads them to propose Knowledge-guided Context Optimization (KgCoOp) where they try to reduce down the distance between hand-crafted prompts and learnable prompts during training.
@@
In our work, we identify some key issues with the aforementioned architectures. First, the features obtained from meta-net in CoCoOp (Zhou et al. 2022a) are global in nature and hence susceptible to issues like clutter and noise in many few-shot and out-of-distribution test cases (we demonstrate this empirically later on). Next, these features are directly added to all the learned prompt vectors, thus resulting in an equal weighting for each prompt. Consequently, this model is unable to learn which of the prompt vectors are more semantically relevant and contextually meaningful during inference, which makes model less generalizable on unseen classes in a zero-shot settings. While Yao, Zhang, and Xu (2023) in KgCoOp architecture prevents the catastrophic knowledge forgetting by minimizing the distance between hand-crafted prompts and learnable prompts, there is no clear evidence of what are the efficient discrete prompts. Moreover, it is hard to design generalized hand-crafted discrete prompts and sometimes tend to overfit linguistic features (Liu et al. 2021).
@@
To address the aforementioned problems, we propose a new technique called Contextual Prompt Learning (CoPL). Our key ideas include aligning prompts to local image context, realized with local features, and determining which prompts are more sematically relevant conditioned on such local context. Our insight is that by doing so, we are able to learn a more appropriate weighting of the prompts, unlike equal weighting of prior work above, that is semantically reflective of the actual content of the image under consideration. Inspired by the concept of global attention from the NLP literature (Luong, Pham, and Manning 2015), during training, we propose to align each local feature vector (e.g., computed from a local image patch) to a set of dynamic soft-prompts using a learned context vector that attends to these prompt vectors. An overview of the methdology is shown in Figure 1. This produces a set of attention weights for the prompt vectors that are semantically aligned to local image regions. This results in CoPL learning being more generalizable features as we demonstrate with an extensive set of zero-shot and few-shot classification evaluation settings.
@@
We conduct a comprehensive set of experiments on visual classification on $11$ different datasets and scenarios (zero-shot, one-shot, seen/unseen, and within-dataset and cross-dataset). Across all these experiments, we demonstrate substantial performance improvement when compared to all the baselines, indicating the ability of our method to be adapted across various classification settings with little or no training and much reduced prompt engineering.
@@
Our key contributions are: (i) We identify two key shortcomings with existing prompt-based image classification methods: equal weighting of the prompt vectors and no flow of contextual local information of input images to the prompt vectors while learning during back-propagation; (ii) We propose CoPL: Contextualized Prompt Learning, a new method that addresses the issues above by learning prompt weights dynamically and aligning the resulting prompt vectors with local image features; (iii) We conduct extensive experiments with CoPL under a variety of classification scenarios and demonstrate substantial performance improvements, particularly in various unseen and few-shot data scenarios. Importantly, on $11$ different image recognition dataset, on average CoPL achieves state-of-the-art performance on unseen class classifications beating state-of-the-art zero-shot large-scale model CLIP by $1.4\%$, conditional prompting model CoCoOp by $3.9\%$, knowledge pruning model KgCoOp by $2.0\%$ and gradient learning method ProGrad by $4.9\%$ on accuracy. We evaluated CoPL on cross-dataset zero-shot image recognition tasks and on an average on $8$ datasets, CoPL outperformed CoCoOp by $2.3\%$ on accuracy. Our extensive experiments on domain generalization in Table 3, established CoPL as state-of-the-art most generalized model by outperforming MaPLe (Khattak et al. 2023) and CoPrompt (Roy and Etemad 2023).
Related Work
Multimodal Models:
@@
Vision-language models have shown great potential in learning generic visual representations. The core idea has been to use natural language supervision for image representation learning and align them jointly in the same embedding space (Jia et al. 2021; Radford et al. 2021). Early explorations in this line of research include related problem formulations - metric learning (Frome et al. 2013a), multilabel classification (Gomez et al. 2017; Joulin et al. 2016), n-gram language learning (Li et al. 2017), captioning (Desai and Johnson 2021). Traditionally, hand-crafted descriptors (Elhoseiny, Saleh, and Elgammal 2013; Socher et al. 2013) were the mode of capturing image representations. Later, convolutional neural net based architectures (Frome et al. 2013b; Ba et al. 2015) were introduced. Recent works have focused on learning joint representations of both the modalities using deep learning architectures (Fürst et al. 2021; Jia et al. 2021; Li et al. 2022b; Kamath et al. 2021). With the introduction of transformers (Vaswani et al. 2017), Li et al., (Li et al. 2019) proposed Visual-BERT, where texts and images are jointly encoded in a single transformers architecture. One of the biggest milestone in the multimodal research is the CLIP model (Radford et al. 2021). It is a dual encoder based model and during training it matches pairs of images and texts. One of the main component of CLIP is the carefully designed prompts which most of the times very hard to formulate. To overcome this, Zhou et al., (Zhou et al. 2022b) designed CoOp, which trains dynamic soft-prompts during back-propagation.
Prompting:
The intuition behind prompt learning is to capture user intention and instructions to perform certain downstream tasks (Li et al. 2022c; Parmar et al. 2022; Zhu et al. 2022). With the introduction of GPT (Brown et al. 2020), prompt engineering is shown to be performing efficiently in few-shot knowledge adaptation. But, building prompt templates is hard and requires immense skills. Recently, researchers have proposed “soft-prompting”. The main intuition is to learn dynamic continuous prompt tokens during back-propagation (Gao, Fisch, and Chen 2021; Jiang et al. 2020; Lester, Al-Rfou, and Constant 2021; Li and Liang 2021; Liu et al. 2021, 2022). Recently, Goswami et al. (Goswami et al. 2023) highlights that the soft prompts can be further tuned with the semantic knowledge of language models without explicitly verbalizers setup. At the same Sarkar et al. (2023) showed tuning these soft prompts can help downstream tasks. CoOp (Zhou et al. 2022b) is designed to train these soft prompts during training. Later Zhou et al. (2022a) introduced CoCoOp on top of CoOp to improve the performance by conditioning on image input. On the other hand, KgCoOp (Yao, Zhang, and Xu 2023) and ProGrad (Zhu et al. 2023) were proposed to improve the performance of CoOp by aligning prompts towards general knowledge. While ProGrad tries to optimize the prompts with the aligned direction, KgCoOp minimises the distance between hand-crafted prompts and learnable prompts during training. As our methodology is directly improving the CoCoOp by infusing the local context of the image during prompt training, we take CoCoOp as our direct baseline. Along with CoOp and CoCoOp, KgCoOp and ProGrad methodologies are directly in line with our way of learning representations, we considered these methods as our standard baselines. Recently, MaPLe (Khattak et al. 2023) and CoPrompt (Roy and Etemad 2023) introduces a new way of learning by infusing prompts to both text and image encoder. Such infusion in a multi-modal setup may affect generalization,and hence in our generalization experiment in Table 3, we compare against these approaches too.
@@
Here, we introduce and discuss details of our proposed method CoPL: Contextualized Prompt Learning. Since CoPL uses the same architectural backbone as CLIP (Radford et al. 2021), we first begin with a brief review followed by a discussion on how our closest baseline, CoCoOp (Zhou et al. 2022a), is trained. While we use the CLIP backbone for simplicity, we are in no way limited by this design choice. In fact, our method is very much applicable to be used in conjunction with a variety of other architectures, e.g., VisualBERT (Li et al. 2019), MDETR (Kamath et al. 2021), GLIP (Li et al. 2022a) etc.
Review of CLIP
@@
CLIP (Radford et al. 2021) is trained using the standard contrastive learning setup where there is a text encoder and an image encoder and the overall objective function is to get their outputs as close as possible in the joint space. The text encoder is implemented with the Transformer model (Vaswani et al. 2017) which takes word sequences as input and produces both individual sequence-level as well as overall sentence-level representations. The image encoder is implemented with the ViT architecture (Dosovitskiy et al. 2021) which produces local (at the patch level) as well as global image features. The contrastive loss function is designed to capture the similarity between the relevant text and images, that is, the cosine similarity between the related text and image will be maximized whereas the cosine similarity with all other unrelated pairs will be minimized. To be precise, for a $K$-class classification problem, the discrete prompts are designed to have one “class” token. The training objective is then to fill the token with the $i^{th}$ class name where $w_{i}$ is the weight vector generated by the text-encoder for the same.
Review of CoOp/CoCoOp
@@
CoCoOp: Conditional Context Optimization is built on top of a previously introduced Context Optimization (CoOp) (Zhou et al. 2022b) algorithm. It is observed that the choice of prompts plays a major role in vision-language understanding (Radford et al. 2021), though it is quite hard to find the best match between prompts and image descriptions. To overcome such a need for prompt engineering, the CoOp model (Zhou et al. 2022b) trains a set of continuous vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}\}$ as context tokens during back-propagation. The CoCoOp (Zhou et al. 2022a) algorithm improves CoOp’s (Zhou et al. 2022b) performance by learning to generate prompts conditioned on each image instance, i.e., these now change for each image unlike in CoOp, where they are fixed. CoCoOp does this by training an additional two-layer neural network called meta-net that takes an image feature vector as input and produces a conditional vector that is combined with the prompt vectors to generate the final image-dependent prompts. This is done as follows:
$\mathbf{v}_{m}(\mathbf{x})=\mathbf{v}_{m}+h_{\theta}(\mathbf{x})$ (1)
@@
Methodology
Figure 2: The overall architecture of CoPL: Contextual Prompt Learning is illustrated in this figure. It improves generalization of pretrained CLIP encoders by using contextual information from image patches, for prompt learning.
CoPL: Contextual Prompt Learning
While CoCoOp generalizes better than CoOp for unseen-class classification, there is significant scope for improvement. First, since CoCoOp uses global feature vectors for learning the updated prompts, it focuses less on the discriminative regions in images (which tend to be more local). Next, the addition operation performed in Equation 1 does not capture the individual importance of each prompt token. This is particularly important since certain discriminative regions in images may weight certain prompts more when compared to others and this is not captured in the CoCoOp model. To address these issues, we propose CoPL, a simple and intuitive algorithm that operates at the local feature granularity while also aligning them with prompts to learn prompt importance weights.
Given an image $\mathbf{I}$, we first compute a set of local feature vectors. For instance, this can be the output of a vision transformer model that generates patch embeddings. Let $\mathbf{s}\in\mathbb{R}^{P\times B\times d}$ (where $P$ is number of patches from image, $B$ is training batch size, and $d$ is feature dimesionality) be this. Conditioned on these local features, we determine the semantically most meaningful prompts. To do this, we take inspiration from the attention work of Luong et al. (Luong, Pham, and Manning 2015) and generate context representations that explicitly consider both the learnable prompt tokens as well as the patch representations. We first learn a lightweight neural network to generate a conditional token for the representation of each patch in $\mathbf{s}$ as:
@@
$=\frac{\text{exp}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i}))}{\sum_{i=1}^{M}(\text{exp}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i})))}$
@@
$\mathbf{v}_{m}(\mathbf{x})=\mathbf{v}_{m}+\sum_{i=1}^{P}\mathbf{c}_{i}$ (6)
In a nutshell, CoPL calculates the prompts for the $i$-th class as $t_{i}=[v_{1}(x),v_{2}(x),...,v_{M}(x),cl_{i}]$, where $cl_{i}$ is the embedding of the $i^{th}$ class label. The prediction probability is calculated as:
@@
where $g()$ is the feature vector produced by the text encoder and $\gamma$ is a temperature parameter. During our entire pipeline, the pre-trained CLIP model is fixed.
Experimental Evaluation
Methodology Protocols Caltech101 OxfordPets StandfordCars Flowers102 Food101 FGVCAircraft SUN397 DTD ImageNet EuroSAT UFC101 Average
CLIP Seen 96.8 91.1 62.3 72.0 90.1 27.1 69.3 53.2 72.4 56.4 70.5 69.3
Unseen 94.0 97.2 74.8 77.8 91.2 36.2 75.3 59.9 68.l 64.0 77.5 74.2
HM 95.4 94.1 68.6 74.8 90.6 31.0 72.2 56.3 70.2 60.0 73.8 71.7
CoOp Seen 98.0 93.6 78.1 97.6 88.3 40.4 80.6 79.4 76.4 92.1 84.6 82.7
Unseen 89.8 95.2 60.4 59.6 82.2 22.3 65.8 41.1 67.8 54.7 56.0 63.2
HM 93.7 94.4 68.1 74.0 85.1 28.7 72.5 54.2 71.9 68.9 67.4 71.6
CoCoOp Seen 97.9 95.2 70.4 94.8 90.7 33.4 79.7 77.0 75.9 87.4 82.3 80.5
Unseen 93.8 97.6 73.4 71.7 91.2 23.7 76.8 56.0 70.4 60.0 73.4 71.7
HM 95.8 96.4 72.0 81.7 90.9 27.7 78.2 64.8 73.1 71.2 77.6 75.8
KgCoOp Seen 97.7 94.6 71.7 95.0 90.5 36.2 80.2 77.5 75.8 85.6 82.8 80.7
Unseen 94.3 97.7 75.0 74.7 91.7 33.5 76.5 54.9 69.9 64.3 76.6 73.6
HM 96.0 96.1 73.3 83.6 91.0 34.8 78.3 64.3 72.7 73.4 79.6 77.0
ProGrad Seen 98.0 95.0 77.6 95.5 90.3 40.5 81.2 77.3 77.0 90.1 84.3 82.4
Unseen 93.8 97.6 68.6 71.8 89.5 27.5 74.1 52.3 66.6 60.8 74.9 70.7
HM 95.9 96.3 72.8 82.0 89.9 32.8 77.5 62.4 71.4 72.6 79.3 76.1
CoPL Seen 98.1 95.6 70.7 96.1 90.9 36.1 80.2 78.0 77.8 89.2 83.1 81.4
Unseen 94.9 97.8 74.4 72.1 91.4 31.3 77.3 50.0 71.3 34.2 76.6 75.6
HM 96.5 96.7 72.6 84.1 91.1 33.7 78.7 64.0 74.5 61.7 79.8 78.5
Table 1: Performance of CoPL and the baselines on $11$ classification datasets; each training dataset consists $16$-shots per class. These results highlights that CoPL has better generalization than other state-of-the-art methods. The results for the base methods are borrowed from KgCoOp (Yao, Zhang, and Xu 2023). HM refers to Harmonic Mean between the seen and unseen classes.